# MMSR: Symbolic Regression is a Multimodal Task

### v4

@@ 3 Method
3.1 Architecture
As depicted in Figure 2, and Algorithm 1 the MMSR consists of two transformer-based encoders, each tailored for learning the symbolic or data representations of mathematical functions. Except that, a decoder receives the encoded numeric representations to decode the features to mathematical expressions. Additionally, an alignment module is designed to match the feature representations of numeric inputs and symbolic inputs. The contrastive learning (Chuang et al., 2020) is used to make the distance between the data and expression skeleton features that are derived from the same expression closer. During training, MMSR receives synthetically created symbolic equations and their associated numeric data as inputs to the symbolic encoder and numeric encoder, respectively. In total, MMSR is trained on approximately 10 million synthetic paired (numeric, symbolic) examples.
@@ 3.1.1 Data Encoder
Considering the permutation invariance of the data feature: it should maintain the same feature with the order change of data input. We thus applied SetTransformer (Lee et al., 2019) to be our data encoder. The input to the encoder consists of the data points $\mathcal{D}=\{X,y\}\in\mathbb{R}^{n\times d}$, which are first passed through a trainable affine layer to project them into a latent space $h_{n}\in\mathbb{R}^{d_{h}}$. The resulting vectors are then passed through several induced set attention blocks (Lee et al., 2019), which are several cross-attention layers. First, cross-attention uses a set of trainable vectors as the queries and the input features as keys and values. Its output is used as the keys and values for the second cross attention, and the original input vectors are used as the queries. After these cross-attention layers, we add a dropout layer (Srivastava et al., 2014). In the end, we compute cross attention between a set of trainable vectors (queries) to fix the size of the output such that it does not depend on the number of input points.
@@ 3.1.2 Expression Skeleton Encoder
Symbolic expression by [+,,*$[+,-,*,/,sin,cos,exp,log\\ sqre,C-1,C-2,...,C4,C5,x_{1},x_{2},...]$ and other basic operators, where [C-5,…,C5] represents a constant placeholder. Each expression can be represented by an expression binary tree, and we can get a symbol sequence by expanding the expression binary tree in preorder, We call this the expression skeleton. Furthermore, we treat each symbol as a $token$ and then embed it.
In particular, to solve the problem that expressions may exhibit high token-level similarities but are suboptimal for equation-specific objectives such as fitting accuracy, which is caused by only using the cross-entropy loss between the generated sequence and the real sequence, we encode the expression symbol sequence and the constant sequence separately. For example, for the expression $12.20.021*sin(x)+$ 12.2, we would have the skeleton [+,*$[+,*,sin,x,C-1,C2]$. And constant sequence $[0,0,0,0.21,0.122]$. It is worth noting that we replace constants with special symbols. The constants are encoded using a scientific-like notation where a constant C is represented as a tuple of the exponent $c_{e}$ and the mantissa $c_{m}$, $C=c_{m}10^{c_{e}}$. Where $c_{e}$ belongs to integers between [-5,5] and $c_{m}$ belongs to [-1,1]. For example, for the constant 0.021 above, we use the placeholder C-1 in the skeleton, where $c_{m}$ is equal to 0.21 and $c_{e}$ is equal to -1, so $0.021=0.21*10^{-1}$.
Figure 3: From the diagram above, we can intuitively see the difference between the one-step strategy and the two-step strategy. The one-step strategy trains the three modules together, while the two-step strategy trains the two encoders and then the decoder.
@@ 3.1.3 Expression Decoder
Algorithm 1 MMSR pre-training
  Input: data: $[X,y]$; Expression Skeleton: $S$; Constants: C; Epoch: $E$; Number of batch: $N$; $\lambda_{1}$; $\lambda_{2}$; $\lambda_{3}$.
  repeat
     Initialize $MMSR=M_{\theta}$.
     for $i=1$ to $E$ do
        for $j=1$ to $N$ do
           $[X_{j},y_{j}]$, [$S_{j}$] = GetBatch($[X,y]$, [$S$])
           $F_{Data}$ = SetTransformer([$X_{j},y_{j}$])
           $F_{Skeleton}$ = SkeletonEncoder([$S_{j}$])
           $S_{pred}$, $C_{pred}$ = ExpressionEncoder($F_{Data}$, $F_{Skeleton}$)
           $\mathcal{L_{CON}}$ = Contrastiveloss($F_{Data}$, $F_{Skeleton}$)
           $\mathcal{L_{CE}}$ = CrossEntropy($S_{i}$,$S_{pred}$)
           $\mathcal{L_{MSE}}$ = MSE($C_{i}$,$C_{pred}$)
           $\mathcal{L}$ = $\lambda_{1}\mathcal{L_{CE}}$+$\lambda_{2}\mathcal{L_{MSE}}$ + $\lambda_{3}\mathcal{L_{CON}}$
        end for
        Compute the gradient $\nabla_{\theta}\mathcal{L}$ and use it to update $\theta$.
     end for
  until Arrival of termination condition
3.1.4 Alignment Module
In the multimodal field, to make the features mapped by the matching modes close, to facilitate the subsequent multimodal feature fusion, to achieve better results. Contrastive loss is introduced in mainstream multimodal models to make the matched data features and skeleton features closer to each other, and the mismatched ones far away from each other in the feature space. Specifically, we align the features of the data obtained by the data point feature extraction extractor with the skeleton features obtained by the skeleton feature extractor. For the data point features, we will get the features of $\mathbb{R}^{b\times k\times h}$ for each batch, where $b$ stands for batch size and $k\times h$ represents the feature size obtained for each data. To facilitate the comparison loss later, we will map the features obtained from each batch from $\mathbb{R}^{b\times k\times h}$ to $\mathbb{R}^{b\times h}$. Similarly, we perform a similar operation for the skeleton feature. Then, the comparison loss is calculated for multiple data and skeleton samples contained in a Batch, and the similarity between positive samples (paired data and skeleton) is expected to be as high as possible, and the similarity between negative samples (unpaired data and skeleton) is expected to be as low as possible. The contrastive loss is formulated as follows.
@@ 3.1.4 Alignment Module
Figure 4: As we can see from the figure, under the premise that other configurations are completely the same, before the introduction of contrastive learning (Figure a), the cosine distance between the matched data features and the skeleton features extracted by the model in the feature space is not significantly smaller than that of the unmatched ones (the values on the diagonal are not significantly smaller than the others). However, after introducing contrastive learning (Figure b), we can clearly see that the values on the diagonal are significantly smaller than the values at other positions. That’s what we expected.
Figure 5: From the above five plots, we can see that as the size of the training data increases, the performance of the model on each dataset continues to improve. When the data size reaches 10M, the model has achieved satisfactory results on the selected data sets, and the average $R^{2}$ on each data set is greater than 0.98. The remaining 8 kinds of data sets are shown in Appendix Figure 7 8
3.2 Training Objective
The loss function plays a crucial role in directing the training process of deep neural networks, which is imperative for effective deep learning. Our chosen loss function is composed of three components: cross-entropy loss ($L_{CE}$), mean squared error(MSE) loss ( $L_{MSE}$ ), and contrastive loss ( $L_{CON}$ ).
@@ 3.2 Training Objective
Here, $\mathcal{L_{CE}}$ represents the cross-entropy loss between the predicted expression sequence and the true sequence. $\mathcal{L_{MSE}}$ represents the MSE loss of the true constants and the predicted constants. $\mathcal{L_{CON}}$ represents the contrastion loss over a batch of samples, which is used to make the data features and expression features that match in the feature space closer to each other, and vice versa. $\lambda_{1}$ $\lambda_{2}$ $\lambda_{3}$ are hyperparameters that allow us to adjust the proportion of the three losses.
Table 1: The effect of the contrastive loss on the performance of the MMSR. For expression 2, let $\lambda_{1}=\lambda_{2}=1$; $\lambda_{3}$ is equal to 0.0. 0.1 and 1.0.
Benchmark Contrastive loss ($\mathcal{L_{CON}}$)

### v5

@@ 3 Method
3.1 Expression Representation
We represent the expression in the form of a binary tree. For example, for the expression $y=4.2*x_{1}+sin(x_{2})$, the form of the expression binary tree is shown in Fig. 2, and we expand it in preorder traversal order to obtain a symbol sequence (expression skeleton). It can be thought of as a ‘sentence’ where each symbol is considered a word. So we can cleverly use generative models to sequentially generate expression skeletons, just like large language models generate sentences.
Figure 2: Expression binary tree. A binary tree representation of the expression $y=4.2*x_{2}+\sin(x_{1})$ and a preorder traversal (skeleton) of its binary tree.
Figure 3: The framework of MMSR model. Among them, the fig. (b) shows the overall algorithm flowchart of MMSR. The fig. (a) shows the structure of the skeleton encoder of MMSR, and the fig. (c) shows the structure of the decoder, note that in the decoder, we use the cross-attention mechanism in the Multi-head cross-attention module for information fusion.
3.2 Architecture of MMSR
As depicted in Fig. 3, and Algorithm 1 the MMSR consists of two transformer-based encoders, each tailored for learning the symbolic or data representations of mathematical functions. Except that, a decoder receives the encoded numeric representations to decode the features to mathematical expressions. Additionally, an alignment module is designed to match the feature representations of numeric inputs and symbolic inputs. The contrastive learning [24] is used to make the distance between the data and expression skeleton features that are derived from the same expression closer. During training, MMSR receives synthetically created symbolic equations and their associated numeric data as inputs to the symbolic encoder and numeric encoder, respectively. In total, MMSR is trained on approximately 10 million synthetic paired (numeric, symbolic) examples.
Input: data: $[X,y]$, Expression Skeleton: $S$, Constants: $C$, Epoch: $E$, Number of batch: $N$, $\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$.
1 repeat
2       Initialize $MMSR=M_{\theta}$.
3       for $i=1$ to $E$ do
4             for $j=1$ to $N$ do
5                   $[X_{j},y_{j}]$, [$S_{j}$] = GetBatch($[X,y]$, [$S$])
6                   $F_{Data}=\text{SetTransformer}([X_{j},y_{j}])$
7                   $F_{Skeleton}=\text{SkeletonEncoder}([S_{j}])$
8                   $S_{pred},C_{pred}=\text{ExpressionDecoder}(F_{Data},F_{Skeleton})$
9                   $\mathcal{L_{CON}}=\text{ContrastiveLoss}(F_{Data},F_{Skeleton})$
10                   $\mathcal{L_{CE}}=\text{CrossEntropy}(S_{i},S_{pred})$
11                   $\mathcal{L_{MSE}}=\text{MSE}(C_{i},C_{pred})$
12                   $\mathcal{L}=\lambda_{1}\mathcal{L_{CE}}+\lambda_{2}\mathcal{L_{MSE}}+\lambda_{3}\mathcal{L_{CON}}$
13
14            Compute the gradient $\nabla_{\theta}\mathcal{L}$ and use it to update $\theta$.
15
16
17until Arrival of termination condition;
Algorithm 1 MMSR
@@ 3.2.1 Data Encoder
Considering the permutation invariance of the data feature: it should maintain the same feature with the order change of data input. We thus applied SetTransformer [23] to be our data encoder. The input to the encoder consists of the data points $\mathcal{D}=\{X,y\}\in\mathbb{R}^{n\times d}$, which are first passed through a trainable affine layer to project them into a latent space $h_{n}\in\mathbb{R}^{d_{h}}$. The resulting vectors are then passed through several induced set attention blocks [23], which are several cross-attention layers. First, cross-attention uses a set of trainable vectors as the queries and the input features as keys and values. Its output is used as the keys and values for the second cross attention, and the original input vectors are used as the queries. After these cross-attention layers, we add a dropout layer [64]. In the end, we compute cross attention between a set of trainable vectors (queries) to fix the size of the output such that it does not depend on the number of input points.
@@ 3.2.2 Expression Skeleton Encoder
Symbolic expression by [+, -, $\times$, $\div$, sin, cos, exp, log
sqre, C-1, C-2,…,C4, C5, $x_{1}$, $x_{2}$,… ] and other basic operators, where [C-5,…,C5] represents a constant placeholder. Each expression can be represented by an expression binary tree, and we can get a symbol sequence by expanding the expression binary tree in preorder, We call this the expression skeleton. Furthermore, we treat each symbol as a $token$ and then embed it.
In particular, to solve the problem that expressions may exhibit high token-level similarities but are suboptimal for equation-specific objectives such as fitting accuracy, which is caused by only using the cross-entropy loss between the generated sequence and the real sequence, we encode the expression symbol sequence and the constant sequence separately. For example, for the expression $12.20.021*sin(x)+$ 12.2, we would have the skeleton [+, $\times$, sin, x, C-1, C2]. And constant sequence $[0,0,0,0.21,0.122]$. It is worth noting that we replace constants with special symbols. The constants are encoded using a scientific-like notation where a constant C is represented as a tuple of the exponent $c_{e}$ and the mantissa $c_{m}$, $C=c_{m}10^{c_{e}}$. Where $c_{e}$ belongs to integers between [-5,5] and $c_{m}$ belongs to [-1,-0.1] $\vee$ [0.1,1]. For example, for the constant 0.021 above, we use the placeholder C-1 in the skeleton, where $c_{m}$ is equal to 0.21 and $c_{e}$ is equal to -1, so $0.021=0.21*10^{-1}$.
@@
Figure 4: The feature fusion process of MMSR model. The fused features are calculated from the feature information embedded implicitly in the data and expression skeleton.
@@ 3.2.3 Expression Decoder
The expression decoder module is primarily responsible for two critical tasks. The first is feature fusion, where aligned data features are combined with skeleton features. The second task is expression generation, which uses these fused features to produce the predicted expression.
  • 1.
    Feature Fusion
Feature fusion is a crucial step in MMSR, as the fused features contain significant information embedded within the data and the expression skeleton. This comprehensive feature enhances the decoder stage, enabling it to generate more accurate and refined target expressions. Specifically, we apply a cross-attention mechanism to achieve this feature fusion, and the fused features are obtained by the following calculation as Eq1.
$&\mathcal{FF}(Q,K,V)&=&\ \text{softmax}\left(\frac{Q\times K^{T}}{\sqrt{d_{k}}}\right)V\\ &\quad\quad Q&=&F_{s}W_{Q}\\ &\quad\quad K&=&F_{xy}W_{K}\\ &\quad\quad V&=&F_{xy}W_{V}\end{aligned}\right.$ (1)
Here, $\mathcal{FF}$ denotes the fused features. $F_{xy}$ represents the data feature, while $F_{S}$ refers to the skeleton feature. $W_{Q}$, $W_{K}$, and $W_{V}$ are mapping matrices responsible for adjusting the dimensions of the corresponding features. $Q$(query) represents the mapped skeleton feature, $K$(keys) can be understood as the description of the data feature, and $V$(value) corresponds to the data feature after matrix mapping. $d_{k}$ serves as the scaling factor.
The detailed process of feature fusion is illustrated in Fig. 4.
First, the data $[X,y]$ is passed through the SetTransformer encoder to obtain the data feature $F_{xy}$.This feature $F_{xy}$ is then multiplied by $W_{v}$ and $W_{k}$ to produce the value vector $V$ and the key vector $K$, respectively. While, the skeleton encoder extracts the skeleton feature $F_{s}$, which is subsequently multiplied by $W_{q}$ to generate the query vector $Q$.
Next, a dot product operation is performed between the query vector $Q$ and the keys vector $K$. Then, through a Softmax function, the attention weight is obtained.
Finally, the attention weight is multiplied by the value vector $V$. So far, we obtained the fused feature that incorporates the data information and the expression skeleton information.
  • 1.
    Skeleton Generation
@@
3.2.4 Modal Alignment Mechanism
In the multi-modal field, modal alignment is usually introduced to make the features mapped by the matching modes close, to facilitate the subsequent multi-modal feature fusion, to achieve better results. In order to achieve the modal alignment. Contrastive loss is introduced in mainstream multi-modal models to make the matched data features and skeleton features closer to each other, and the mismatched ones far away from each other in the feature space. Specifically, we align the features of the data obtained by the data point feature extraction extractor with the skeleton features obtained by the skeleton feature extractor. For the data point features, we will get the features of $\mathbb{R}^{b\times k\times h}$ for each batch, where $b$ stands for batch size and $k\times h$ represents the feature size obtained for each data. To facilitate the comparison loss later, we will map the features obtained from each batch from $\mathbb{R}^{b\times k\times h}$ to $\mathbb{R}^{b\times h}$. Similarly, we perform a similar operation for the skeleton feature. Then, the comparison loss is calculated for multiple data and skeleton samples contained in a Batch, and the similarity between positive samples (paired data and skeleton) is expected to be as high as possible, and the similarity between negative samples (unpaired data and skeleton) is expected to be as low as possible. The contrastive loss is formulated as follows.
@@ 3.2.4 Modal Alignment Mechanism
where $x_{i}$ and $y_{j}$ are embeddings of the data in the $i$-th pair and that of the sequence of expression in the $i$-th pair. $N$ is the batch size, and $\theta$ is the temperature to scale the logits.
Futhermore, positive and negative samples are constructed for contrastive learning. Positive samples represent expressions that match with specific data, while negative samples refer to expressions that do not match the given data. The aim of contrastive learning is to enable the model to distinguish between expressions that match (positive samples) and do not match (negative samples) with the data. This recognition capacity is achieved through continual optimization and learning throughout the model’s training process. In MMSR, the data encoder is employed to extract features from the data, and the expression skeleton encoder is used to extract features from the expression skeletons, aiming to bring the matching data and expression skeletons as close together as possible in the feature space, whereas ensuring that the mismatches are kept as far apart as possible. By doing so, alignment across different modalities is achieved. This process aids in better learning and understanding of the associations between data features and expression features, enhancing the performance of multimodal solutions to symbolic regression tasks.
3.3 Joint Training Strategy
We adopt one-step training strategy to train the network, as shown in Fig. 3. The loss function plays a crucial role in directing the training process of deep neural networks, which is imperative for effective deep learning. We design a multi-loss function consisting of three components: cross-entropy loss ($\mathcal{L_{CE}}$), mean squared error(MSE) loss ( $\mathcal{L_{MSE}}$ ), and contrastive loss ( $\mathcal{L_{CON}}$ ).
@@ 3.3 Joint Training Strategy
Here, $\mathcal{L_{CE}}$ represents the cross-entropy loss between the predicted expression sequence and the true sequence:
$\mathcal{L_{CE}}=-\sum_{k}^{M}\log P_{MMSR}(\hat{s}_{k}|\hat{s}_{1:k-1},X,Y)$ (4)
where M is the sequence length and $\hat{s}_{k}$ is the $k$-th predicted symbol, $\hat{s}_{1:k-1}$ denotes the first $k$-1 symbols in the predicted sequence.
$\mathcal{L_{MSE}}$ represents the MSE loss of the true constants and the predicted constants:
$\mathcal{L_{MSE}}=\frac{1}{M}\sum_{k}^{M}\frac{(c_{k}-\hat{c}_{k})^{2}}{(c-\overline{c})^{2}}$ (5)
Here, M denotes the length of the predicted constant sequence, $c_{k}$ denotes the $k$-th true value in the constant sequence, and $\hat{c}_{k}$ denotes the $k$-th value in the predicted constant sequence. $\overline{c}$ denotes the mean of the real constant sequence.
@@
$\mathcal{L_{CON}}$ represents the contraction loss over a batch of samples, which is used to make the data features and expression features that match in the feature space closer to each other, and vice versa. The specific formula is given in Eq. 2. In MMSR, we achieve modal alignment by introducing contrastive loss $\mathcal{L_{CON}}$ into the loss function.
$\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$ are hyperparameters that allow us to adjust the proportion of the three losses.
4 Experiments
4.1 Dataset
  • 1.
    Training dataset: The magnitude of the training data is 10 million (M), which is generated based on the expression skeleton. After obtaining the expression skeleton by random sampling, the constant encoding is generated and the values [-1,-0.1$\cup$[0.1,1] are randomly assigned to the locations where the constant placeholders [C-5,…,C5] occurrence in the expression skeleton. Then, we can get an expression $y=f(X)$. For each expression, we collect a random set of data $[X,y]$. Finally, we generated 10M expressions and gained 10M sets of data for MMSR training. For instance, the expression skeleton [sin, $\times$, C1, $x$] is obtained by random sampling and then the constant encoding [0.0, 0.0, 0.23, 0.0] is generated. This yields the encoding of the expression y=sin(2.3$\times$$x$). Then, with a set of $x$-values, a corresponding set of $y$-values can be computed. In this way, we obtained a set of training data. For the sake of a fair comparison, the magnitude of training data for the baseline methods is also set at 10M, which were generated in a similar manner according to the format requirements of each respective method.
  • 2.
    Testing dataset: In order to test the performance of MMSR and baselines, we conducted comparative analyses using prevalent datasets, designated as ’Nguyen’, ’Keijzer’, ’Korns’, ’Constant’, ’Livermore’, ’Vladislavleva’, ’R’,’ Jin’, ’Neat’,’ Others’, ’Feynman’, ’Strogatz’ and ’Black-box’. The three datasets Feynman ’, ’Strogatz’ and ’Black-box’ are from the SRbetch dataset. These datasets contain nearly 400 expressions in total. We believe that they can accurately and comprehensively test the performance of each algorithm.
Table 1: The results of performance comparison. At a 0.95 confidence level, a comparison of the coefficient of determination ($R^{2}$) and the expression complexity(Nodes) conducted between MMSR and four baselines.
Group Dataset MMSR TPSR[25] End2End[56] NeSymReS[19] SymbolicGPT[18]
$R^{2}$ Nodes $R^{2}$ Nodes $R^{2}$ Nodes $R^{2}$ Nodes $R^{2}$ Nodes
Standards Nguyen $\textbf{0.9999}_{\pm 0.001}$ 14.5 $0.9948_{\pm 0.002}$ 16.0 $0.8814_{\pm 0.004}$ 16.3 $0.8568_{\pm 0.003}$ 18.2 $0.6713_{\pm 0.005}$ 21.6
Keijzer $\textbf{0.9983}_{\pm 0.003}$ 16.3 $0.9828_{\pm 0.003}$ 20.6 $0.8134_{\pm 0.005}$ 18.4 $0.7992_{\pm 0.003}$ 21.3 $0.6031_{\pm 0.004}$ 24.5
Korns $\textbf{0.9982}_{\pm 0.003}$ 19.2 $0.9325_{\pm 0.004}$ 22.9 $0.8715_{\pm 0.004}$ 23.4 $0.8011_{\pm 0.005}$ 24.1 $0.6613_{\pm 0.005}$ 29.2
Constant $\textbf{0.9986}_{\pm 0.002}$ 24.5 $0.9319_{\pm 0.002}$ 35.3 $0.8015_{\pm 0.003}$ 28.3 $0.8344_{\pm 0.003}$ 32.9 $0.7024_{\pm 0.004}$ 38.5
Livermore $\textbf{0.9844}_{\pm 0.003}$ 29.4 $0.8820_{\pm 0.004}$ 38.2 $0.7015_{\pm 0.004}$ 32.2 $0.6836_{\pm 0.005}$ 36.2 $0.5631_{\pm 0.0005}$ 41.2
Vladislavleva $\textbf{0.9862}_{\pm 0.003}$ 21.7 $0.9028_{\pm 0.005}$ 24.6 $0.7422_{\pm 0.005}$ 22.2 $0.6892_{\pm 0.004}$ 27.3 $0.5413_{\pm 0.004}$ 36.6
R $\textbf{0.9924}_{\pm 0.004}$ 16.4 $0.9422_{\pm 0.003}$ 16.2 $0.8512_{\pm 0.004}$ 19.5 $0.7703_{\pm 0.005}$ 19.9 $0.7042_{\pm 0.005}$ 25.2
Jin $\textbf{0.9943}_{\pm 0.003}$ 28.3 $0.9826_{\pm 0.004}$ 29.5 $0.8611_{\pm 0.004}$ 29.8 $0.8327_{\pm 0.003}$ 32.2 $0.7724_{\pm 0.006}$ 36.9
Neat $\textbf{0.9972}_{\pm 0.004}$ 17.3 $0.9319_{\pm 0.002}$ 16.4 $0.8044_{\pm 0.004}$ 19.7 $0.7596_{\pm 0.005}$ 20.6 $0.6377_{\pm 0.005}$ 26.4
Others $\textbf{0.9988}_{\pm 0.002}$ 20.6 $0.9667_{\pm 0.002}$ 22.5 $0.8415_{\pm 0.003}$ 22.3 $0.8026_{\pm 0.003}$ 23.5 $0.7031_{\pm 0.004}$ 31.8
SRBench Feynman $\textbf{0.9913}_{\pm 0.002}$ 20.8 $0.8928_{\pm 0.004}$ 21.3 $0.7353_{\pm 0.004}$ 22.0 $0.7025_{\pm 0.005}$ 22.4 $0.5377_{\pm 0.005}$ 26.8
Strogatz $\textbf{0.9819}_{\pm 0.003}$ 21.6 $0.8249_{\pm 0.002}$ 24.4 $0.6626_{\pm 0.003}$ 25.4 $0.6022_{\pm 0.003}$ 28.1 $0.5229_{\pm 0.004}$ 32.6
Black-box $\textbf{0.9937}_{\pm 0.004}$ 26.7 $0.8753_{\pm 0.004}$ 29.3 $0.6925_{\pm 0.004}$ 31.2 $0.6525_{\pm 0.005}$ 33.9 $0.5833_{\pm 0.005}$ 37.4
Average 0.9934 21.3 $0.9264$ 24.4 $0.7892$ 23.9 $0.7528$ 26.2 $0.6311$ 32.4
4.2 Compared Methods
  • 1.
    TPSR[25]: One uses a pre-trained model as a policy network to guide the MCTS process, which greatly improves the search efficiency of MCTS.
  • 2.
    End2End[56]: End2End further encodes the constants so that the model generates the constants directly when generating the expression, abandoning the constant placeholder ‘C’ used by the previous two algorithms.
  • 3.
    NeSymReS[19]: Based on SymbolicGPT, the algorithm treated each ‘operator’ (e.g. ‘sin’, ‘cos’) as a token. The sequence of expressions is generated in turn. The setup is more reasonable.
  • 4.
    SymbolicGPT[18]: The classical algorithm for pre-trained models, which treats the symbolic regression problem as a translation problem. Treat a ‘single letter’ (e.g. ‘s’, ‘i’, ‘n’) as a token. Generating expressions sequentially.
4.3 Evaluation Indicators
4.3.1 Fitting Ability
Coefficient of determination ($R^{2}$) is utilized as the metric to evaluate the performance. Formally, $R^{2}$ is calculated as follows:
$R^{2}=1-\sum_{i=1}^{n}\frac{(y_{i}-\hat{y_{i}})^{2}}{(y_{i}-\overline{y})^{2}}$ (6)
Where $n$ is the number of sample points, $y_{i}$ is the $i^{th}$ true value, $\hat{y_{i}}$ is the $i^{th}$ predicted value, and $\overline{y}$ is the mean of the true y values.
4.3.2 Expression complexity
Symbolic regression aims to find mathematical expressions that are concise and explanatory. Therefore, when evaluating the performance of different symbolic regression algorithms, the complexity of the expressions generated by the algorithms always be considered.
Therefore, in this work, we evaluate the complexity of each method by counting the length of the preorder traversal of the final expression produced by each method, i.e., the number of symbols. For example in Fig. 2 expression $y=4.2*x_{2}+\ sin(x_{1})$, the first sequence traversal of expression binary tree is [+,*, c1, $x_{2}$, sin, $x_{1}$], a total of 6 nodes (symbols), so the expression of complexity is 6.
4.3.3 Model Complexity
We measure the complexity of a model in terms of the number of model parameters and time complexity (inference time).
  • 1.
    Number of parameters. The number of all parameters in the model that need to be optimized updated.
  • 2.
    Time complexity. The amount of time, in seconds, from the input to the final result.
Figure 5: Radar charts demonstrating the fitting ability ($R^{2}$). It can be found that MMSR achieved the best fitting performance among all the comparison methods.
4.4 Results
4.4.1 Performance Evaluation of MMSR
In Regression Symbolic, $R^{2}$ is the most basic and important metric to reflect the fitting ability of the algorithm. In the paper, each expression in the test set is tested 20 times, and then the average of its $R^{2}$ is taken as the final result, with a confidence interval of 0.95. The performance comparison between our algorithm and four baselines is shown in Table 1 and visualised with radar chart in Fig.5 intuitively. From the experiment results, we can find that MMSR achieved the best fitting performance among all the comparison methods.
Furthermore, we run 20 times for each test sample of all datasets and subsequently compute the average complexity of the resulting expressions. Detailed statistical results regarding the average complexity of each dataset are listed in Table 1. The expressions obtained through the proposed MMSR method have the lowest average complexity. This makes the expressions extracted from the data using the MMSR method easier to interpret compared to other methods. This is the goal that MMSR is dedicated to, obtaining more concise and more accurate expressions.
Figure 6: Distance of data feature pairs after normalization. (a) without modal alignment, (b) with modal alignment. As we can see from the figure, under the premise that other configurations are completely the same, before the introduction of contrastive learning (a), the cosine distance between the matched data features and the skeleton features extracted by the model in the feature space is not significantly smaller than that of the unmatched ones (the values on the diagonal are not significantly smaller than the others). However, after introducing contrastive learning (b), we can clearly see that the values on the diagonal are significantly smaller than the values at other positions. That’s what we expected.
4.4.2 Effect of Modal Alignment on Features
To verify whether the extracted features before and after the introduction of contrastive learning are as expected: Data and skeleton features that match each other are close to each other in the feature space, whereas those that do not match are far away from each other. We first randomly generate 50 expression samples, feed these samples into the model without contrastive learning, and then obtain 50 pairs of features for the data and skeleton. We calculate the cosine distance between the extracted 50 data features and 50 skeleton features and then obtain the 50*50 cosine distance matrix A. Finally, we normalize the matrix A and visualize it to get Fig. 6(a). Similarly, we feed these samples into a model that uses contrastive learning, resulting in heatmap 6(b).
From Fig. 6, we can see that there is indeed a large change in the extracted features before and after the introduction of contrastive learning. In Fig. 6(a), we can see that the data on the diagonal is not significantly smaller than the other positions, indicating that the matching data features and skeleton features are not the closest to each other in the feature space. On the contrary, we can clearly find in Figure 3b that the data on the diagonal is significantly smaller than the other positions, which indicates that our contrastive learning plays a role. The features of the two modalities, data and skeleton, have been well aligned.
4.4.3 Ablation Study of Modal Alignment
In MMSR, we achieve modal alignment by introducing contrastive loss ($\mathcal{L_{CON}}$) into the loss function ($\mathcal{L}$), and denote the degree of the contribution of modal alignment by $\lambda_{3}$. To investigate the effectiveness of modal alignment to the symbolic regression tasks, we conduct an ablation experiment consisting of three groups. In each group, the settings of the parameters $\theta$, $\lambda_{1}$ and $\lambda_{2}$ were kept consistent, where $\theta$ = 0.07, $\lambda_{1}$ = $\lambda_{2}$ =1. And the parameter $\lambda_{3}$ was set to 0.0, 0.1 and 1.0 respectively. When $\lambda_{3}$ = 0.0, it indicates that modal alignment is not used. When $\lambda_{3}$ = 0.1 or 1.0, it indicates that the modal alignment is applied, and the larger the value of $\lambda_{3}$, the greater the role that modal alignment is plays in the MMSR method.
Table 2: The results of ablation experiments examining the effect of modal alignment.
Dataset $R^{2}$
Hyperparameters $\theta=0.07$ $\theta=0.07$ $\theta=0.07$
$\lambda_{1}=1.0$ $\lambda_{1}=1.0$ $\lambda_{1}=1.0$
$\lambda_{2}=1.0$ $\lambda_{2}=1.0$ $\lambda_{2}=1.0$